# Multi-mode Coherent Detection Ghost Imaging Lidar and Vibration-Mode Imaging

### v1

@@ 3.1 Experiments of static target imaging in the bucket-detector-based multi-mode CD-GI lidar
Moreover, we investigated the statistical characteristics of the IF signal energy in the bucket-detector-based multi-mode CD-GI lidar system and its relationship with the total energy of the reconstructed image pixels. The experimental results in Fig. 4 (2.a)–(2.b) show that the ratio of the total energy of all reconstructed image pixels to the coherent summation of the IF signal energy varies with detector size. Specifically, the measured ratios were 623.4 for $D_{\mathrm{det}}=300\mu m$ and 271.2 for $D_{\mathrm{det}}=200\mu m$. These ratios closely match the theoretical expectations based on the effective number of pixels shown in Eq. (12) (697 and 256, respectively, as detailed in Supplement 1, Sec. S4.A), directly validating the ability to decouple the coherence sum of the reflected multi-mode light fields to individual pixels using field correlation.
These two experimental results jointly demonstrate that, although the IF signal energy is constrained by Siegman antenna theorem[56], the exploitation of field correlation enables the reflected multi-mode light fields collected by the bucket detector to be coherently decoupled into effective single-mode components determined by the pixel size, thereby effectively circumventing the constraint imposed by Siegman antenna theorem.
3.3 Experiments of vibration modes
Building on these validations, the bucket-detector-based multi-mode CD-GI lidar is further applied in experiments to assess its capability in extracting target vibration modes.
3.3.1 Experiments of one vibrating target:
@@
A loudspeaker (8 $\Omega$, 5 W) was employed as the vibration source. Because the diaphragm material of the speaker has very low reflectivity at 1550 nm, a high-reflectivity (HR) film was affixed to its surface to enhance the optical reflectivity. The target reflectivity distribution was set as Fig. 5 (a). The loudspeaker was driven by a sinusoidal voltage signal generated from a computer sound card, with vibration frequencies set to 1000 Hz and 2000 Hz. The DMD operated at an inversion repetition frequency of 1000 Hz.
@@
Fig. 5 (b) and (c) display the vibration characteristics of the target at different excitation frequencies, including the time frequency spectrum of the vibration target, the reconstructed instantaneous coherent image at different time slices, and the superposed intensity image. Since the reconstructed vibration height are dynamic, they are provided as videos in the Supplementary Materials (Visualizations 1.1 and 1.2). These results demonstrate that the vibration mode of the target can be successfully reconstructed.
To investigate the changes in vibration modes at different driving frequencies, the target was configured as a strip-shaped object measuring 7 mm in width and 60 mm in length. The HR film was clamped at both ends, while the central region was adhered to a diaphragm using adhesive tape, forming a planar vibrating target. A simplified schematic of the experimental setup is shown in Fig. 6 (a). The vibration frequencies were set to 1000 Hz, 1500 Hz, and 2000 Hz. These frequencies were selected to assess the system’s capability to detect and reconstruct micro-vibration characteristics across different temporal scales.
@@
Figure 6: Simplified schematic of the single vibrating target setup: (a), the vibrating feature information fusion diagram for one independent vibrating target, and each row corresponds to a different vibration frequency: (b)–(d) represent 1000 Hz, 1500 Hz, and 2000 Hz, respectively, from top to bottom. The coherent image shows the instantaneous intensity reflectivity image of the target reconstructed via field correlation. Then, the vibration feature image presents the vibration amplitude of the target at different vibration frequencies across uniformly spaced time slices. The last column shows the reconstructed longitudinal cross-sectional image of the vibration. The dark blue curve is the outer contour line of the vibration amplitude superimposed on each other.
@@
3.3.2 The Imaging Experiment for Distinguishing Multiple Vibration Modes
To further evaluate the system’s capability to simultaneously resolve multiple vibrating targets, a dual-target experiment was conducted. In this configuration, two strip-shaped vibrating targets with identical dimensions (7 mm in width and 60 mm in length) were positioned at different axial distances and independently driven at 500 Hz and 1000 Hz, respectively. The corresponding target distances were 65.00 m and 65.68 m. A simplified schematic of the experimental setup is shown in Fig. 7.
Figure 7: Multidimensional information reconstruction results for dual vibrating targets: time-frequency information, complex reflectivity information, vibration mode information and distance information.(For the dynamic video, see Visualization 3). (a) Schematic of the experimental setup with transmitting and receiving optical paths. (b) Received time-domain waveform from the balanced bucket detector. (c) STFT processing and image reconstruction for different time slices. (d) Reconstructed vibration mode images corresponding to 500 Hz and 1000 Hz targets. (e) Coherent detection results showing the fusion information. (f) FFT processing of the time-domain waveform signal for range extraction. (g) Reconstructed range images for two targets at 72.34 m and 73.10 m. (h) Intensity detection results of the targets.
Fig. 7 presents the multidimensional information retrieved from the dual vibrating targets. First, feature extraction is performed on the time frequency spectrum obtained via STFT processing on Fig. 7 (c). By employing a curve-fitting approach, the vibration frequencies of the two targets are identified as 495 Hz and 1007 Hz. Next, based on temporal slicing, the vibration amplitude corresponding to each frequency component is calculated as shown in Fig. 7 (d), enabling the separation and reconstruction of the time-varying vibration mode maps of the two targets with distinct vibration frequencies. The dynamic evolution of the target vibration modes is provided in Visualization 3.
Then, pulse compression was applied to the electrical signal, which involves performing a FFT over the entire signal to extract the range information of the two targets as shown in Fig. 7 (f) and (g). The reconstructed target distances were 72.34 m (500 Hz) and 73.10 m (1000 Hz), yielding a longitudinal separation of 0.766 m. Compared with the actual separation of 0.680 m, the resulting deviation of 0.087 m falls well within the system’s range resolution of $\Delta z=c/(2B)\approx 0.187$ m, indicating reliable range discrimination. After determining the distance of each target, combining the range information with the reconstructed vibration modes allows for more precise target characterization.
Finally, Fig. 7 (e) and (h) compare the target information reconstructed via intensity correlation and field correlation. Compared with intensity image, the proposed system enables simultaneous acquisition of target reflectivity, vibration frequency discrimination, and three-dimensional imaging.
4 Conclusion and Discussion
In summary, this work presents a CD-GI lidar strategy based on bucket detector of reflected multi-mode light fields. The system improves detection efficiency by employing a bucket detector. Meanwhile, the bucket-detector-based multi-mode CD-GI lidar system breaks the limitations of Siegman antenna theorem by exploiting field correlation to decouple the coherent summation of reflected multi-mode light fields. The results demonstrate that this strategy reconstructs the static spatial distribution of targets with high sensitivity under the spatiotemporal bandwidth product of a single detector and simultaneously retrieves their amplitude and phase information, enabling complex reflectivity imaging. In addition, this strategy not only discriminates targets according to their vibration frequencies, but also reconstructs the spatial distribution of targets’ vibration modes. Moreover, the three-dimensional imaging capability is preserved, enabling a more comprehensive characterization of the target’s spatiotemporal characteristics.
This work shows potential in applications such as autonomous sensing, non-contact structural health monitoring, precision micro-vibration measurement, and multi-modal remote sensing. By simultaneously acquiring time–frequency features and spatial vibration mode distributions, the system provides enriched target descriptors that enable more accurate recognition, classification, and comprehensive target analysis.
@@ 4 Conclusion and Discussion
Funding :
[…]
Acknowledgment The authors would like to thank Professor Wenlin Gong for con-structive discussions.

### v2

@@ 3.3.1 Experiments of one vibrating target:
The data processing workflow is implemented as follows:
  1. 1.
    Time-Frequency Representation: The electrical signals corresponding to 8,192 patterns are processed via STFT using a rectangular window ($l=4096$, $d=2048$).
  2. 2.
    Turbulence Compensation: To mitigate phase jitter induced by atmospheric turbulence, the resulting STFT spectra are processed according to Eq. (8). Specifically, the phase fluctuations are compensated by performing point-wise multiplication between the STFT spectra of odd frames and the complex conjugate of the even frames.
  3. 3.
    Micro-Doppler Extraction: The modules of the 4,096 compensated spectra are accumulated and normalized. A global thresholding algorithm is then applied to the normalized spectrogram to identify the support set $\omega_{k}^{(h)}$, which represents the time-frequency locations of all micro-Doppler components.
  4. 4.
    Coordinate and Height Imaging:
    • Step 1: The identified micro-Doppler frequencies are converted into a vibration height matrix $A_{k}^{(h)}$ using Eq. (10).
    • Step 2: A looping algorithm extracts the complex signals at $f=\frac{\omega_{k}^{(h)}}{2\pi}$ from each compensated STFT spectrum to serve as the bucket signals.
  5. 5.
    Field Correlation Reconstruction: For each temporal slice $k$, the bucket signals are correlated with the reference fields. This process reconstructs the complex reflectivity of the target at specific micro-Doppler frequencies. The spatial coordinates are determined by extracting the intensity of the reconstructed images.
  6. 6.
    Spatiotemporal Synthesis: By integrating the vibration height matrix with the corresponding spatial coordinates across all temporal slices, the full-field spatial distribution of the target’s vibration mode is successfully retrieved.
@@
Fig. 5 (b) and (c) display the vibration characteristics of the target at different excitation frequencies, including the time frequency spectrum, the reconstructed amplitude image at different time slices, and the intensity image obtained by summing the intensity images from different time slices. Since the reconstructed vibration height images are dynamic, they are provided as videos in the Supplementary Materials (Visualizations 1.1 and 1.2). These results demonstrate that the vibration mode of the target can be successfully reconstructed.
@@
Figure 6: Simplified schematic of the single vibrating target setup: (a), the vibrating feature information fusion diagram for one vibrating target, and each row corresponds to a different vibration frequency: (b)–(d) represent 1000 Hz, 1500 Hz, and 2000 Hz, respectively, from top to bottom. First column: Time-frequency distribution profiles characterizing the target vibration. Second column: Reconstructed amplitude images of the target’s complex reflectivity obtained through field correlation. Third column: Corresponding vibration height images reconstructed at different time slices.
To investigate the changes in vibration modes at different driving frequencies, the target was configured as a strip-shaped object measuring 7 mm in width and 60 mm in length. The HR film was clamped at both ends, while the central region was adhered to a diaphragm using adhesive tape, forming a planar vibrating target. The effective coding region is $64\times 64$ modulation array after 4-pixels binning (Hadamard patterns). The detection FoV is measured at $77$ mm. The repetition frequency of the DMD is set at $1000$ Hz. A simplified schematic of the experimental setup is shown in Fig. 6 (a). The vibration frequencies were set at 1000 Hz, 1500 Hz, and 2000 Hz.
According to the vibration modes imaging reconstruction approach for vibrating targets, Fig. 6 (b)-(d) illustrates the reconstructed vibrating feature information of the target at different vibration frequencies. Under a fixed driving power, the vibration height of target gradually decreases with increasing vibration frequency[58], which aligns with the physical behavior of a damped driven harmonic oscillator. Since the reconstructed vibration images are dynamic, the complete temporal evolutions are provided as Supplementary Media (Visualizations 2.1, 2.2, and 2.3). For static demonstration in the manuscript, representative frames are sampled from the sequences with an initial offset at the 7th frame and a step size of 15 frames.
3.3.2 The Imaging Experiment for Distinguishing Multiple Vibration targets
To further evaluate the system’s capability to simultaneously resolve multiple vibrating targets, two strip-shaped vibrating targets with identical dimensions (7 mm in width and 60 mm in length) were positioned at different axial distances and independently driven at 500 Hz and 1000 Hz, respectively. The corresponding target distances were 65.00 m and 65.68 m. The effective coding region is $64\times 64$ modulation array after 4-pixels binning (Hadamard patterns). The repetition frequency of the DMD is set at $1000$ Hz. The detection FoV is measured at $77$ mm. A simplified schematic of the experimental setup is shown in Fig. 7 (a).
Figure 7: Results of the vibration frequencies, vibration mode spatial distributions, 2D spatial distributions, and distances of two targets with different vibration frequencies (For the dynamic video, see Visualization 3). (a) Schematic of the experimental setup. (b) Received time-domain waveform from the balanced bucket detector (Taking the electrical signal corresponding to a single pattern as an example). (c) STFT processing. (d) Reconstructed vibration height images at different time slices corresponding to 1000 Hz and 500 Hz targets. (e) Reconstructed intensity images of the target’s complex reflectivity. (f) FFT processing for distance extraction. (g) Reconstructed depth images for two targets at 65.14 m and 65.90 m. (h) The reconstructed spatial distribution of two targets’ vibration modes.
Fig. 7 presents the results of the vibration frequencies, vibration mode spatial distributions, 2D spatial distributions, and distances of two targets with different vibration frequencies. The raw time-domain waveform captured by the balanced bucket detector is presented in Fig. 7 (b). The corresponding time-frequency spectrum, obtained by performing STFT on the time-domain waveform, is shown in Fig. 7 (c). First, feature extraction is performed on this spectrum. By employing a curve-fitting approach, the vibration frequencies of the two targets are identified as 495 Hz and 1007 Hz. Next, based on temporal slicing, the vibration height corresponding to each frequency component is calculated as shown in Fig. 7 (d), enabling the separation and reconstruction of the time-varying vibration height images of the two targets with distinct vibration frequencies. The dynamic evolution of the target vibration modes is provided in Visualization 3. Fig. 7 (e) represent the reconstructed intensity image of the targets’ complex reflectivity.
Finally, pulse compression could also be applied to the electrical signal, which involves performing a FFT over the entire signal to extract the distance information of the two targets as shown in Fig. 7 (f) and (g). The reconstructed target distances were 65.14 m (500 Hz) and 65.90 m (1000 Hz), yielding a longitudinal separation of 0.76 m. Compared with the actual separation of 0.68 m, the resulting deviation of 0.08 m falls well within the system’s distance resolution of $\Delta z=c/(2B)\approx 0.187$ m, indicating reliable distance discrimination. Combining the distance information with the reconstructed vibration modes allows for more precise target characterization. Compared with intensity image, our system enables simultaneous acquisition of spatial distribution of vibration modes and target reflectivity.
4 Conclusion and Discussion
In summary, this work presents a multi-mode CD-GI lidar strategy based on bucket detector. It breaks the limitations of Siegman antenna theorem by exploiting field correlation to decouple the coherent summation of reflected multi-mode light fields. The results demonstrate that this strategy reconstructs the spatial distribution of target’s complex reflectivity with high sensitivity under the spatiotemporal bandwidth product of a single bucket detector. This strategy not only discriminates targets according to their vibration frequencies, but also reconstructs the spatial distribution of targets’ vibration modes. Moreover, the depth imaging and three-dimensional imaging capability is preserved, enabling a more comprehensive characterization of the target’s spatiotemporal characteristics.
By simultaneously acquiring time–frequency features and spatial vibration mode distributions, the system provides enriched target descriptors that enable more accurate recognition, classification, and comprehensive target analysis. This work shows potential in applications such as autonomous sensing, non-contact structural health monitoring, precision micro-vibration measurement, and multi-modal remote sensing.
@@ 4 Conclusion and Discussion
Funding This work is supported by the National Key R&D Program of China (JCKY2024110C034).
@@
Acknowledgment The authors are grateful to Professor Wenlin Gong for his constructive discussions.